What is meant by (a.) polarization by reflection and (b.) angle of polarization?

    Ca!calculate the angle of polarization for water of refractive index 1.33. 


1
Expert's answer
2021-02-24T12:50:40-0500

Answer

angle between the reflected and refracted ray is called the Brewster angle or polarization angle.

Angle of polarization is given by the "\theta=tan^{-1}(\mu)" .



If refrective idex with respect to water is 1.33 then ange of polarization will be

"\theta=tan^{-1}(1.33) \\\theta=53.06°"